# Whistle: Data-Efficient Multilingual and Crosslingual Speech Recognition via Weakly Phonetic Supervision

### v1

@@ Abstract
There exist three approaches for multilingual and crosslingual automatic speech recognition (MCL-ASR) - supervised pre-training with phonetic or graphemic transcription, and self-supervised pre-training. We find that pre-training with phonetic supervision has been underappreciated so far for MCL-ASR, while conceptually it is more advantageous for information sharing between different languages. This paper explores the approach of pre-training with weakly phonetic supervision towards data-efficient MCL-ASR, which is called Whistle. We relax the requirement of gold-standard human-validated phonetic transcripts, and obtain International Phonetic Alphabet (IPA) based transcription by leveraging the LanguageNet grapheme-to-phoneme (G2P) models. We construct a common experimental setup based on the CommonVoice dataset, called CV-Lang10, with 10 seen languages and 2 unseen languages. A set of experiments are conducted on CV-Lang10 to compare, as fair as possible, the three approaches under the common setup for MCL-ASR. Experiments demonstrate the advantages of phoneme-based models (Whistle) for MCL-ASR, in terms of speech recognition for seen languages, crosslingual performance for unseen languages with different amounts of few-shot data, overcoming catastrophic forgetting, and training efficiency. It is found that when training data is more limited, phoneme supervision can achieve better results compared to subword supervision and self-supervision, thereby providing higher data-efficiency. To support reproducibility and promote future research along this direction, we will release the code, models and data for the whole pipeline of Whistle at https://github.com/thu-spmi/CAT upon publication.
@@ I Introduction
In multilingual speech recognition, training data for a number of languages, often referred to as seen languages, are merged to train a multilingual model, which can be used to recognize speech from all seen languages. The multilingual model can also serve as a pre-trained model, which can be further fine-tuned for crosslingual speech recognition. Crosslingual speech recognition refers to recognizing utterances in a new language, which is unseen in training the multilingual model. From machine learning perspective, such multilingual and crosslingual training can be regarded as performing multi-task learning and transfer learning, which promotes sharing of statistical strength. The advantage is that the ASR performance for low-resource languages, both seen and unseen, can be improved, and the cost of system building and maintenance for multiple languages can be reduced as well.
@@
The general concept of multilingual and crosslingual speech recognition has been applied for a long time, dating back to the time when GMM-HMM based classic models and then DNN-HMM based hybrid models are prevalent in ASR research, to name a few, e.g., in [2] and [3] respectively. Recently, end-to-end models have emerged [4, 5, 6], which can be directly trained from phonetic or graphemic transcription, eliminating the first pass of producing HMM state alignment. For end-to-end models, the approach of pre-training followed by fine-tuning has attracted increasing interests and achieved good performance. There are mainly two classes of pre-training methods, based on either self-supervised learning or supervised learning. Self-supervised pre-training is conducted over unlabeled speech data from multiple languages for speech representation learning in general [7, 8, 9]. Supervised pre-training, by applying end-to-end models on multilingual labeled speech data, can be further divided into two sub-categories of research, which are contrasted by using different types of modeling units. The first is grapheme-based or subword-based [10, 11, 12, 13], which, collectively referred to as based on graphemic transcription (orthography), creates a shared token set across multiple languages, e.g., using 10K sentence pieces [11]. The second trains end-to-end models on phonetic transcriptions [14, 15, 16, 17, 18], which usually utilizes International Phonetic Alphabet (IPA) symbols to create a (nearly-)universal phone inventory, e.g., using 187 phones [14].
@@
Intuitively, the key to successful multilingual and crosslingual recognition is to optimize information sharing during multilingual training and maximize the knowledge transferring from a well trained multilingual model to the model trained for recognizing utterances in a new language [15]. Taking this perspective, we could examine the pros and cons of the three approaches - supervised pre-training with graphemic transcription or phonetic transcription, and self-supervised pre-training, which is detailed in Section II.
@@
While requiring pronunciation lexicons, pre-training with phonetic supervision is more advantageous for information sharing between different languages. For phonetic supervision, IPA symbols include enough symbols to represent the fundamental sounds of all languages, and sounds in different languages share these phonetic representations [19]. In contrast, graphemes and subwords are in fact from writing systems of languages (orthography), not for describing and distinguishing all the sounds in human language throughout the world, which is exactly phonetic transcription does. Creating a graphemic token set from multiple languages for supervision is non-trivial and delicately affects ASR performance; until recently, tokenization strategy is still under investigation and needs a balance between granularity and ASR performance [12]; adding new languages for crosslingual recognition further complicates the design of tokenization. Besides the above theoretical analysis of supervised pre-training with graphemic transcription and phonetic transcription, an interesting research question is about empirical comparison. It has been empirically found that compared to learning with graphemic supervision, learning with phonetic supervision performs equally strong and tends to be more data-efficient in monolingual ASR [20, 21, 22, 23]. But to the best of our knowledge, there have been no solid experiments to study which approach is better or if they yields similar results for MCL-ASR, when evaluated in a common experimental setup (Research Question 1, referred to as RQ-1).
To address the problem of requiring phonetic transcription for phonetic supervision, we note that phonetic resources and tools have been steadily developed over these years and are easily accessible, including grapheme-to-phoneme (G2P) models and tools [24, 25, 26], phoneme inventories [27]. We can relax the requirement of gold-standard human-validated transcripts, and in this paper, we obtain the IPA phonetic transcripts by leveraging the LanguageNet G2P models [25]. The LanguageNet G2P models are available for 142 languages, with the phoneme error rates (PERs) ranging from 7% to 45%. So the main aim of this paper is to investigate weakly supervised pre-training with somewhat noisy phonetic transcription. This is in spirit similar to the work in Whisper [13]. But instead of using weakly graphemic supervision in Whisper, our work employs weakly phonetic supervision. We call the approach investigated in this paper: Whistle (Weakly phonetic supervision strategy for multilingual and crosslingual speech recognition).
@@
A secondary interesting research question is to compare supervised pre-training and self-supervised/un-supervised pre-training. Basically, we agree with the comments in [13]. Current pre-trained models for speech such as based on wav2vec 2.0 [28] aim to learn speech representation in general over unlabeled data; They mostly are encoder-only and thus lack an equivalently performant decoder, which requires at least adding a classifier layer and supervised finetuning over labeled data even for seen languages. These comments, presumably, are suited to comparing self-supervision to both graphemic supervision [13] and phonetic supervision (our work). These being said, to the best of our knowledge, there have been no strict experiments to study which approach is better or if they yields similar results for MCL-ASR, when evaluated in equal settings (Research Question 2, referred to as RQ-2).
In summary, this paper explores supervised pre-training with weakly phonetic supervision, towards data-efficient multilingual and crosslingual speech recognition. Our main contributions are as follows.
@@
  • We construct a common experimental setup based on the CommonVoice dataset, called CV-Lang10, to evaluate multilingual and crosslingual speech recognition, with 10 seen languages and 2 unseen languages, measuring both phoneme error rate (PER) and word error rate (WER). A set of experiments are conducted on CV-Lang10 to compare, as fair as possible, the three approaches under the common setup - supervised pre-training with graphemic transcription or weakly phonetic transcription, and self-supervised pre-training for MCL-ASR. These experiments present our effort to answer RQ-1 and RQ-2.
@@
  • Many prior works on multilingual and crosslingual speech recognition were conducted on internal or proprietary datasets such as GlobalPhone [29] and IARPA Babel [^1] https://www.iarpa.gov/index.php/research-programs/babel, which are not openly-available. We find that supervised pre-training with phonetic supervision has been underappreciated so far for MCL-ASR. To promote future research along this direction, we release the code, models and data for the whole pipeline of Whistle at the following URL: https://github.com/thu-spmi/CAT.
@@ II-A MCL-ASR with phonetic supervision
Research in multilingual and cross-lingual ASR has long been motivated by phonetics and has used phonetic supervision, e.g., in [30, 3, 2, 14, 31, 15, 17], to name a few. The major phonetic alphbet in use is the International Phonetic Alphabet (IPA), which includes modified Roman letters and diacritics, by means of which the sounds of all human languages can be represented [19]. So a common practice is to combine the phonetic inventory of all languages to be recognized into a global phoneme set, often based on IPA. Employing phonetic units is, presumably, the most intuitive way to promote information sharing and learn language-universal representations for MCL-ASR. Modeling based on phonetic supervision further allows to pursue finer level of information sharing by decomposing phones into a list of phonological articulatory attributes [32, 15, 17, 33].
@@ II-B MCL-ASR with graphemic supervision
Thanks to larger and larger amounts of transcribed speech data and increasingly large neural networks, subword-based supervised pre-training has obtained better and better performance and become a widely adopted strategy in industry to build MCL-ASR systems for increasingly many languages. For example, the Whisper [13] models use the a Byte-Pair Encoding (BPE) text tokenizer and are trained over 680,000 hours cleaned web data by weakly graphemic supervision, capable of recognizing speech from 97 languages. While achieving impressive performance, recent advances in large MCL-ASR models are presumably an effect of scaling power, and it is hard to argue that the good results are not due to having additional data, nor due to the large neural architecture. It remains unclear which approach (phonetic supervision or grapheme supervision) is better when evaluated in an equal experimental setting, or if they produce similar results for MCL-ASR. This paper presents our preliminary effort to answer this question (RQ-1).
@@ II-C MCL-ASR with self-supervision
Self-supervised learning methods mainly refer to some recent learning methods based on contrastive learning such as wav2vec 2.0 [28] or masking prediction such as BERT [37], which can still be regarded as unsupervised learning methods from a classical perspective (no data annotation is required). Therefore, the literature often does not strictly distinguish between unsupervised and self-supervised learning methods in terms of terminology, and we can collectively referred to them as unsupervised learning methods. Self-supervised learning methods such as wav2vec 2.0 [28] have been proposed to learn speech representation in general from multilingual unlabeled speech data. Based on wav2vec 2.0, XLS-R models [8] are trained on unlabeled data from 128 languages. In the recent Massively Multilingual Speech (MMS) project [9], wav2vec 2.0 based models are pre-trained over 1,406 languages, and CTC based multilingual ASR models for 1,107 languages are then fine-tuned using labeled data for each language. Specifically, a linear layer is added on top of pre-trained MMS models which maps to an output vocabulary which is the set of letters in the labeled training data, and is then fine-tuned with the CTC loss.
@@
As commented in [13], while current unsupervised pre-training has improved the quality of audio encoders, the lack of an equivalently high-quality pre-trained decoder is a crucial weakness which limits their usefulness. In the following, we provide a closely related comment. We find that current unsupervised pre-training methods in learning audio encoders such as wav2vec 2.0 does not satisfy the so-called principled unsupervised learning, since “the unsupervised objective may be unrelated to the supervised task of interest” [38]. In contrast, the GPT based unsupervised pre-training method for natural language processing (NLP) tasks is principled, since the supervised objective is the same as (closely related to) the unsupervised objective but only evaluated on a subset of the sequence in NLP [39]. For ASR tasks, these comments favor supervised pre-training (either grapheme-supervision or phonetic supervision) over the current unsupervised pre-training. These being said, remarkably, it has been known in various machine learning tasks that supervised and unsupervised training methods are not mutually exclusive and could be jointly used to define semi-supervised learning, e.g., in image classification [40], speech recognition [41, 42], natural language labeling [43], dialog systems [44]. A complete investigation into semi-supervised learning for ASR is outside the scope of this paper. This paper presents a straightforward empirical comparison between self-supervision and phonetic supervision for MCL-ASR in a common experimental setup (RQ-2).
@@
Figure 1: Illustration of the pre-training and fine-tuning procedures with phonetic supervision, subword supervision, and self-supervision.

### v2

@@ Abstract
There exist three approaches for multilingual and crosslingual automatic speech recognition (MCL-ASR) - supervised pretraining with phonetic or graphemic transcription, and self-supervised pretraining. We find that pretraining with phonetic supervision has been underappreciated so far for MCL-ASR, while conceptually it is more advantageous for information sharing between different languages. This paper explores the approach of pretraining with weakly phonetic supervision towards data-efficient MCL-ASR, which is called Whistle. We relax the requirement of gold-standard human-validated phonetic transcripts, and obtain International Phonetic Alphabet (IPA) based transcription by leveraging the LanguageNet grapheme-to-phoneme (G2P) models. We construct a common experimental setup based on the CommonVoice dataset, called CV-Lang10, with 10 seen languages and 2 unseen languages. A set of experiments are conducted on CV-Lang10 to compare, as fair as possible, the three approaches under the common setup for MCL-ASR. Experiments demonstrate the advantages of phoneme-based models (Whistle) for MCL-ASR, in terms of speech recognition for seen languages, crosslingual performance for unseen languages with different amounts of few-shot data, overcoming catastrophic forgetting, and training efficiency. It is found that when training data is more limited, phoneme supervision can achieve better results compared to subword supervision and self-supervision, thereby providing higher data-efficiency. To support reproducibility and promote future research along this direction, we release the code, models and data for the entire pipeline of Whistle at https://github.com/thu-spmi/CAT/tree/master/egs/cv-lang10.
@@ I Introduction
In multilingual speech recognition, training data for a number of languages, often referred to as seen languages, are merged to train a multilingual model, which can be used to recognize speech from all seen languages. The multilingual model can also serve as a pretrained model, which can be further finetuned for crosslingual speech recognition. Crosslingual speech recognition refers to recognizing utterances in a new language, which is unseen in training the multilingual model. From machine learning perspective, such multilingual and crosslingual training can be regarded as performing multi-task learning and transfer learning, which promotes sharing of statistical strength. The advantage is that the ASR performance for low-resource languages, both seen and unseen, can be improved, and the cost of system building and maintenance for multiple languages can be reduced as well.
@@
The general concept of multilingual and crosslingual speech recognition has been applied for a long time, dating back to the time when GMM-HMM based classic models and then DNN-HMM based hybrid models are prevalent in ASR research, to name a few, e.g., in [2] and [3] respectively. Recently, end-to-end models have emerged [4, 5, 6], which can be directly trained from phonetic or graphemic transcription, eliminating the first pass of producing HMM state alignment as used in DNN-HMM based hybrid models. For end-to-end models, the approach of pretraining followed by finetuning has attracted increasing interests and achieved good performance. There are mainly two classes of pretraining methods, based on either self-supervised learning or supervised learning. Self-supervised pretraining is conducted over unlabeled speech data from multiple languages for speech representation learning in general [7, 8, 9]. Supervised pretraining, by applying end-to-end models on multilingual labeled speech data, can be further divided into two sub-categories of research, which are contrasted by using different types of modeling units. The first is grapheme-based or subword-based [10, 11, 12, 13], which, collectively referred to as based on graphemic transcription (orthography), creates a shared token set across multiple languages, e.g., using 10K sentence pieces [11]. The second trains end-to-end models on phonetic transcriptions [14, 15, 16, 17, 18], which usually utilizes International Phonetic Alphabet (IPA) symbols to create a (nearly-)universal phone inventory, e.g., using 187 phones [14].
@@
Intuitively, the key to successful multilingual and crosslingual recognition is to optimize information sharing during multilingual training and maximize the knowledge transferring from a well trained multilingual model to the model trained for recognizing utterances in a new language [15]. Taking this perspective, we can examine the pros and cons of the three approaches - supervised pretraining with graphemic transcription or phonetic transcription, and self-supervised pretraining, which is detailed in Section II. And there have two interesting research questions (RQs).
@@
The first question is about the comparison between phonetic supervision and graphemic supervision in pretraining for MCL-ASR. While requiring pronunciation lexicons, pretraining with phonetic supervision is more advantageous for information sharing between different languages. For phonetic supervision, IPA symbols include enough symbols to represent the fundamental sounds of all languages, and sounds in different languages share these phonetic representations [19]. In contrast, graphemes and subwords are in fact from writing systems of languages (orthography), not for describing and distinguishing all the sounds in human language throughout the world, which is exactly phonetic transcription does. Creating a graphemic token set from multiple languages for supervision is non-trivial and delicately affects ASR performance; until recently, tokenization strategy is still under investigation and needs a balance between granularity and ASR performance [12]; adding new languages for crosslingual recognition further complicates the design of tokenization. Besides the above theoretical analysis of supervised pretraining with graphemic transcription and phonetic transcription, an interesting research question is about empirical comparison. It has been empirically found that compared to learning with graphemic supervision, learning with phonetic supervision performs equally strong and tends to be more data-efficient in monolingual ASR [20, 21, 22, 23]. But to the best of our knowledge, there have been no solid experiments to study which approach is better or if they yields similar results for MCL-ASR, when evaluated in a common experimental setup (Research Question 1, referred to as RQ-1).
@@
The second interesting research question is to compare supervised pretraining and self-supervised/unsupervised pretraining. Basically, we agree with the comments in [13]. Current pretrained models for speech such as based on wav2vec 2.0 [24] aim to learn speech representation in general over unlabeled data; They mostly are encoder-only and thus lack an equivalently performant decoder, which requires at least adding a classifier layer and supervised finetuning over labeled data even for seen languages. These comments, presumably, are suited to comparing self-supervision to both graphemic supervision [13] and phonetic supervision (our work). These being said, to the best of our knowledge, there have been no strict experiments to study which approach is better or if they yields similar results for MCL-ASR, when evaluated in equal settings (Research Question 2, referred to as RQ-2).
Remarkably, in evaluating and comparing different pretraining approaches to answer the above research questions for MCL-ASR, data-efficiency is an important aspect. Data-efficiency can refer to many things [^1] https://en.wikipedia.org/wiki/Data_efficiency, and in this paper, it mainly refers to achieving better performance with the same amount of data. In particular, data-efficiency in MCL-ASR entails: 1) efficient use of pretraining data (referred to as pretraining data-efficiency or multilingual data-efficiency); and 2) efficient use of limited target language finetuning data such as 1 or 10 hours (referred to as finetuning data-efficiency or crosslingual data-efficiency). The former evaluates the performance of a pretrained, multilingual model in recognizing seen languages, while the latter measures the performance of finetuned models in recognizing unseen languages.
@@
In this paper, we present our effort to answer the above two research questions. Our main contributions are as follows. First, we construct a common experimental setup based on the CommonVoice dataset, called CV-Lang10, to evaluate multilingual and crosslingual speech recognition, with 10 seen languages and 2 unseen languages, measuring both phoneme error rate (PER) and word error rate (WER). A set of experiments are conducted on CV-Lang10 to compare, as fair as possible, the three approaches under the common setup - supervised pretraining with graphemic transcription or phonetic transcription, and self-supervised pretraining for MCL-ASR. It is found in our experiments that phonetic supervision obtains better multilingual data-efficiency than graphemic supervision [^2] Note that a multilingual model from self-supervised pretraining is essentially just an encoder and thus alone cannot be directly applied in recognizing even seen languages. So for the self-supervision approach, we only examine crosslingual data-efficiency and do not consider multilingual data-efficiency. See more discussions at the end of Section II-C.; and compared to both graphemic supervision and self-supervision, phonetic supervision excels in crosslingual data-efficiency.
Second, to address the problem of requiring phonetic transcription for phonetic supervision, we note that phonetic resources and tools have been steadily developed over these years and are easily accessible, including grapheme-to-phoneme (G2P) models and tools [25, 26, 27], phoneme inventories [28]. We can relax the requirement of human-validated gold-standard transcripts, and in this paper we obtain the IPA phonetic transcripts by leveraging the LanguageNet G2P models [26]. The LanguageNet G2P models are available for 142 languages, with the phoneme error rates (PERs) ranging from 7% to 45%. So the main technical aim of this paper is to investigate weakly supervised pretraining with somewhat noisy phonetic transcription. This is in spirit similar to the work in Whisper [13]. But instead of using weakly graphemic supervision in Whisper, our work employs weakly phonetic supervision. We call the approach investigated in this paper: Whistle (Weakly phonetic supervision strategy for multilingual and crosslingual speech recognition).
@@
Third, many prior works on multilingual and crosslingual speech recognition were conducted on internal or proprietary datasets such as GlobalPhone [29] and IARPA Babel [^3] https://www.iarpa.gov/index.php/research-programs/babel, which are not openly-available. We find that supervised pretraining with phonetic supervision has been underappreciated so far for MCL-ASR. To promote future research along this direction, we release the code, models and data for the entire pipeline of Whistle at the following URL: https://github.com/thu-spmi/CAT/tree/master/egs/cv-lang10.
@@ II-A MCL-ASR with phonetic supervision
Research in multilingual and cross-lingual ASR has long been motivated by phonetics and has used phonetic supervision, e.g., in [30, 3, 2, 14, 31, 15, 17], to name a few. The major phonetic alphabet in use is the International Phonetic Alphabet (IPA), which includes modified Roman letters and diacritics, by means of which the sounds of all human languages can be represented [19]. So a common practice is to combine the phonetic inventory of all languages to be recognized into a global phoneme set, often based on IPA. Employing phonetic units is, presumably, the most intuitive way to promote information sharing and learn language-universal representations for MCL-ASR. Modeling based on phonetic supervision further allows to pursue finer level of information sharing by decomposing phones into a list of phonological articulatory attributes [32, 15, 17, 33].
@@ II-B MCL-ASR with graphemic supervision
Thanks to larger and larger amounts of transcribed speech data and increasingly large neural networks, subword-based supervised pretraining has obtained better and better performance and become a widely adopted strategy in industry to build MCL-ASR systems for increasingly many languages. For example, the Whisper [13] models use the a Byte-Pair Encoding (BPE) text tokenizer and are trained over 680,000 hours cleaned web data by weakly graphemic supervision, capable of recognizing speech from 97 languages. While achieving impressive performance, recent advances in large MCL-ASR models are presumably an effect of scaling power, and it is hard to argue that the good results are not due to having additional data, nor due to the large neural architecture. It remains unclear which approach (phonetic supervision or grapheme supervision) is better when evaluated in an equal experimental setting, or if they produce similar results for MCL-ASR. This paper presents our preliminary effort to answer this question (RQ-1).
@@ II-C MCL-ASR with self-supervision
Self-supervised learning methods mainly refer to some recent learning methods based on contrastive learning such as wav2vec 2.0 [24] or masking prediction such as BERT [37], which can still be regarded as unsupervised learning methods from a classical perspective (no data annotation is required). Therefore, the literature often does not strictly distinguish between unsupervised and self-supervised learning methods in terms of terminology, and we can collectively refer to them as unsupervised learning methods. Self-supervised learning methods such as wav2vec 2.0 [24] have been proposed to learn speech representation in general from multilingual unlabeled speech data. Based on wav2vec 2.0, XLS-R models [8] are trained on unlabeled data from 128 languages. In the recent Massively Multilingual Speech (MMS) project [9], wav2vec 2.0 based models are pretrained over 1,406 languages, and CTC based multilingual ASR models for 1,107 languages are then finetuned using labeled data for each language. Specifically, a linear layer is added on top of pretrained MMS models which maps to an output vocabulary which is the set of letters in the labeled training data, and is then finetuned with the CTC loss.
@@
As commented in [13], while current unsupervised pretraining has improved the quality of audio encoders, the lack of an equivalently high quality pretrained decoder is a crucial weakness which limits their usefulness. In the following, we provide a closely related comment. We find that current unsupervised pretraining methods in learning audio encoders such as wav2vec 2.0 does not satisfy the so-called principled unsupervised learning, since “the unsupervised objective may be unrelated to the supervised task of interest” [38]. In contrast, the GPT based unsupervised pretraining method for natural language processing (NLP) tasks is principled, since the supervised objective is the same as (closely related to) the unsupervised objective but only evaluated on a subset of the sequence in NLP [39]. For ASR tasks, these comments favor supervised pretraining (either grapheme-supervision or phonetic supervision) over the current unsupervised pretraining. These being said, remarkably, it has been known in various machine learning tasks that supervised and unsupervised training methods are not mutually exclusive and could be jointly used to define semi-supervised learning, e.g., in image classification [40], speech recognition [41, 42, 43], natural language labeling [44], dialog systems [45]. A complete investigation into semi-supervised learning for ASR is outside the scope of this paper. This paper presents a straightforward empirical comparison between self-supervision and phonetic supervision for MCL-ASR in a common experimental setup (RQ-2).
@@
Figure 1: Illustration of the pretraining and finetuning procedures with (a) phonetic supervision, (b) subword supervision, and (c) self-supervision.
Finally, MCL-ASR with self-supervised, auto-generated phonetic, grapheme or sub-word units based labels is very interesting, which, to our understanding, is closely related to recent progress in unsupervised speech recognition [46, 47, 48], i.e. learning a speech recognizer with only unpaired speech and text. The speech signals are automatically encoded into representative vectors by a pretrained, self-supervised speech encoder such as wav2vec. The representative vectors are then clustered into acoustic tokens, and each speech utterance is represented as a cluster index sequence. It turns out to be very difficult to transcribe the discovered acoustic tokens into phonemes or graphemes in an unsupervised way, though recently there are some progresses, e.g. by GAN [46, 47] or skipgram and positional unigram matching [48]. The dominant way to use pretrained, self-supervised speech encoder is to finetune with phoneme or grapheme labels, as shown in Figure 1(c), which is exactly what we evaluate and compare in the experiments. Investigating self-supervised, auto-generated phonetic or graphemic labels, as a related work in unsupervised ASR, is interesting, but out of the scope of this paper.